# Process-Supervised Reward Models for Verifying Clinical Note Generation: A Scalable Approach Guided by Domain Expertise

### v2

@@ Abstract
Process-supervised reward models (PRMs), which verify large language model (LLM) outputs step-by-step, have achieved significant success in mathematical and coding problems. However, their application to other domains remains largely unexplored. In this work, we train a PRM to provide step-level reward signals for clinical notes generated by LLMs from patient-doctor dialogues. Guided by real-world clinician expertise, we carefully designed step definitions for clinical notes and utilized Gemini-Pro 1.5 to automatically generate process supervision data at scale. Our proposed PRM, trained on the LLaMA-3.1 8B instruct model, outperformed both Gemini-Pro 1.5 and the vanilla outcome-supervised reward model (ORM) in two key evaluations: (1) selecting gold-reference samples from error-containing ones, achieving 98.8% accuracy (versus 70.0% for the vanilla ORM and 93.8% for Gemini-Pro 1.5), and (2) selecting physician-preferred notes, achieving 56.2% accuracy (compared to 37.5% for the vanilla ORM and 50.0% for Gemini-Pro 1.5). Additionally, we conducted ablation studies to determine optimal loss functions and data selection strategies, along with physician reader studies to explore predictors of downstream Best-of-N performance. Our promising results suggest the potential of PRMs to extend beyond the clinical domain, offering a scalable and effective solution for diverse generative tasks.  [^1] Our code is available at https://github.com/hanyin88/prm-clinic.
@@
Hanyin Wang1,2, Chufan Gao2, Qiping Xu1, Bolun Liu1, Guleid Hussein1, Hariprasad Korsapati1, Mohamad El Labban1, Kingsley Iheasirim1, , Mohamed Hassan1, Gokhan Anil1, Brian Bartlett1, Jimeng Sun2 1 Mayo Clinic Health System, 2 University of Illinois Urbana-Champaign wang.hanyin@mayo.edu
@@ 1 Introduction
Figure 1: Overview of study design. We train a PRM to step-by-step verify clinical notes generated by LLMs from patient-doctor dialogues. Step 1: Clinical notes are transformed into a hierarchical structure of steps, designed based on domain expertise to capture the key considerations in clinical documentation. Step 2: Gemini Pro 1.5 is utilized to generate synthetic errors from predefined categories. These errors are systematically swapped with the original steps to create negative samples. Step 3: The PRM is trained using LLaMA-3.1 8B instruct, followed by ablation studies and physician reader studies to evaluate its performance.
@@
Large Language Models (LLMs) show promising potential in generating clinical notes Van Veen et al. (2024), but their outputs can contain errors and may not always align with physician preferences Omiye et al. (2024); Jin et al. (2024). Currently, no automated, scalable method exists to evaluate the quality of LLM-generated clinical notes, making manual clinician evaluation the gold standard. As a result, costly clinician pilot studies are required to assess LLM-based products, such as ambient AI scribes, in real-world healthcare settings Liu et al. (2024).
@@
A potential solution to this challenge is to verify LLM outputs using a reward model (RM). Reward models are integral to Reinforcement Learning from Human Feedback (RLHF) for LLM post-training Ouyang et al. (2022); Stiennon et al. (2020). When employed as verifiers, RMs are categorized into two types: outcome-supervised reward models (ORMs) Cobbe et al. (2021) and, more recently, process-supervised reward models (PRMs) Lightman et al. (2023). ORMs evaluate an entire generation and assign a single reward score at the end, whereas PRMs provide reward scores at each step of the generation, offering several notable advantages. PRMs enable more fine-grained verification of LLM outputs, improving explainability by pinpointing errors at their exact location. This step-wise reward signaling also facilitates advanced inference-time computation techniques, such as Monte Carlo Tree Search (MCTS), and supports step-by-step reinforcement learning Wang et al. (2023b); Snell et al. (2024).
@@
However, despite significant success in mathematical and coding tasks Lightman et al. (2023); Wang et al. (2023b); Ma et al. (2023), the application of PRMs to other domain remains underexplored Beeching et al. . This gap arises from several key challenges: (1) Verification of correctness: Unlike mathematical or coding problems, which have objective ground-truth answers, verifying correctness in open-ended generative tasks is inherently more complex. (2) Process-supervised data collection: Training PRMs requires datasets with annotations for each step of the answer, which makes it challenging to scale when relying on human annotators.
To address these challenges, we propose an effective method for training PRMs to verify open-ended text generation in the clinical domain. We focus on generating clinical notes from patient-doctor dialogues, an emerging application of LLMs in healthcare Barr et al. (2024). Our approach leverages synthetic data to efficiently generate process-supervised datasets at scale by intentionally introducing errors into high-quality notes. Drawing on domain expertise, we design “errors” and structure “steps” to reflect real-world clinical documentation needs. Our developed PRM, trained on the LLaMA-3.1 8B instruct model, outperformed both Gemini-Pro 1.5 and the vanilla ORM in two key evaluations: (1) selecting gold-reference samples from error-containing ones, achieving 98.8% accuracy (versus 70.0% for the vanilla ORM and 93.8% for Gemini-Pro 1.5), and (2) selecting physician-preferred notes, achieving 56.2% accuracy (compared to 37.5% for the vanilla ORM and 50.0% for Gemini-Pro 1.5). The model also exhibited robust performance in out-of-distribution data. Our main contributions are as follows:
@@
    PRMs for verifying clinical note generation: We demonstrate that PRMs can effectively provide step-level reward signals for clinical note generation. In real-world deployment, such PRMs serve two practical purposes: (1) automating the evaluation of AI-generated clinical notes to reduce costly clinician reviews, and (2) guiding LLMs to produce higher-quality notes through inference-time computation.
@@
    Broadening PRM applications beyond math and code: This work expands the application of PRMs beyond their traditional focus on math and coding, demonstrating their potential for open-ended text tasks across diverse domains.
@@
    Scalable process supervision methodology: We propose a methodology that leverages domain expertise to automatically generate process supervision data at scale.
  4. 4.
    New dataset for the research community: We release our process supervision and physician preference dataset to support future research.
@@ 2 Background and Related Work
Recent research highlights a growing interest in process supervision of LLMs to improve their reasoning capabilities. In contrast to earlier work on ORMs Cobbe et al. (2021); Yu et al. (2024), PRMs assess the correctness of LLM outputs step-by-step Lightman et al. (2023); Uesato et al. (2022). One of the core applications of PRMs is Best-of-N sampling, where N complete solutions are generated, and the most suitable one is selected based on the PRM’s reward signal. This method has proven exceptionally effective in mathematical reasoning and code generation tasks, frequently surpassing ORMs Wang et al. (2023b); Ma et al. (2023). Furthermore, PRMs can facilitate step-level reinforcement learning and guide inference-time compute and scaling through techniques such as MCTS Wang et al. (2023b); Ma et al. (2023); Snell et al. (2024). They are considered crucial to the success of advanced reasoning models like OpenAI’s O1. Jiang et al. (2024); Qin et al. (2024); Wang et al. (2024b).
@@
Relevant to our work, the need for human-labeled annotations for each step of PRM training data is prohibitively expensive to scale. To address this limitation, several studies have proposed methods to automatically generate process supervision data, including techniques such as Monte Carlo estimation Wang et al. (2023b); Luo et al. (2024); Wang et al. (2024c). However, these approaches are primarily confined to the mathematical domain, where they are effective due to the unambiguous and easily verifiable nature of mathematical solutions. Our exploration of PRM is also related to prior work on evaluating step-by-step reasoning chains Hao et al. (2024), though our focus is on open-ended text generation tasks, which present unique challenges.
@@
Ambient AI products that transcribe patient-doctor conversations into clinical notes represent one of the most prominent use cases of LLMs in medicine, significantly reducing the burden of clinical documentation Barr et al. (2024). This task was explored at the 2023 ACL ClinicalNLP and CLEFImage workshops, where the best results were achieved using proprietary models Abacha et al. (2023); Yim et al. (2023a). Recent studies have also shown promise in fine-tuning open-source LLMs to generate expert-level clinical notes Wang et al. (2024a, 2023a); Li et al. (2024), though challenges such as inaccuracies and hallucinations remain.
@@ 3 Our Approach: Building PRMs to Verify Clinical Note Generation
In this section, we first introduce the problem and task, followed by an explanation of the core method for constructing the process supervision dataset. We then discuss the training, usage, and evaluation methods for PRMs. An overview of our study is illustrated in Figure 1.
@@ 3.1 Problem Formation
Given a patient-doctor dialogue, an LLM can generate clinical notes comparable to those written by physicians. We aim to train PRMs to effectively verify LLM-generated notes step-by-step, enabling them to evaluate multiple candidates and select the most accurate and physician-preferred note. Guided by recent work (see details in Appendix D.3), we focus exclusively on the most challenging “Assessment and Plan” (A&P) section of the note.
@@ 3.2 Training Data Construction
Unlike mathematical problems, where each solution step is typically a single sentence, the concept of a “step” in an open-ended text generation task is less clear. The A&P sections of clinical notes are semi-structured, consisting of multiple problems, each encompassing several clinical reasoning narratives related to their assessment and plan (Figure 1 and 2). While various design options exist, we adopted a heuristic approach informed by physician expertise, as detailed below.
@@
  • Factual Inaccuracy: Errors involving incorrect information or topics that are referenced in the conversation but not supported by its content.
@@
We tasked Gemini Pro 1.5 with generating a pool of unique errors for each error type in each case. This was achieved through careful prompt engineering and manual inspection of the errors by our physician co-authors. Subsequently, we randomly swapped original steps in the gold-reference notes with entries from the error pools, ensuring that each error was used only once. At this stage, we introduced a fourth error type, Incompleteness, by randomly removing specific steps or entire problems from the samples.
@@
Table 1: Summary statistics of PRM-Clinic. The reported numbers are calculated solely from negative samples. Samples from the same case share the same dialogues. During training, gold-reference samples are mixed with negative samples.
Assign Step Scores
Each step is assigned a score label of “$+$” or “--”. The “$+$” label represents a correct step, and the “--” label indicates an incorrect or incomplete step. For gold-reference samples, every step receives a “$+$” label. For samples containing errors, the erroneous step is assigned a score label of “--”. For samples with incompleteness, we adjust the corresponding problem-level or note-level completeness score label to “--”. Lastly, for any samples containing errors, we assign a score label of “--” to the end-of-note step.
@@
A-Prefer Physician-preference ACI-BENCH tasks 2 and 3 subsets 80 OOD Includes physician preference data used in the first round of RLHF in Wang et al. (2024a). Each case contains three notes generated by an early phase of LLaMA-Clinic.
A-Verify Gold-reference ACI-BENCH tasks 2 and 3 subsets 80 OOD Gold-reference notes generated by an later phase of LLaMA-Clinic. Negative samples are introduced in the same manner as PRM-Clinic.
@@
Figure 3: Main results. PRM outperforms both the vanilla ORM and Gemini Pro models across all tasks. Our model can function as either a PRM or ORM. When operating as a PRM, it outperforms our ORM in the A-Prefer task and performs comparably to the ORM in verification tasks. For A-Prefer, the task involves selecting the physician-preferred note from three candidate notes. For all other tasks, the objective is to identify the correct gold-reference sample from negative samples. For ID tasks, the notes are generated by Gemini Pro, while for OOD tasks, the notes are generated by LLaMA-Clinic. The results of the best-performing PRM are presented alongside the ORM results from the same checkpoint. ID: In-distribution. OOD: Out-of-distribution.
@@ 3.3 Training of PRMs
The training corpus consists of dialogue concatenated with clinical notes, where each step in the notes is separated by a step token followed by a score label token. We used LLaMA-3.1 8B Instruct Dubey et al. (2024) as the base model, utilizing its reserved special tokens for all step and score label tokens.
@@ 3.4 Usage of PRMs
At inference time, we perform a single PRM forward pass over the entire clinical note to obtain PRM’s predicted score label for each step. The clinical note itself is generated by a separate LLM. We define the PRM score as the probability that a given step is correct, which is the softmax probability of the special “$+$” token at that step position.
@@ 3.5 Evaluation of PRMs
We evaluated the reward model’s ability to select the superior note within the Best-of-N framework using two key metrics: (1) Accuracy of selecting the gold-reference note, which measures the model’s ability to identify the gold-reference note among error-containing samples, and (2) Accuracy of selecting the physician-preferred note, which assesses its capability to identify the physician-preferred note from a set of candidate notes.
@@
We compare our model against three baselines: LLM-as-a-judge using Gemini Pro 1.0 and 1.5, as well as a vanilla ORM implementation. In the vanilla ORM, the training corpus lacks step-level annotations and includes only a single score label for the entire clinical note at the end. For ablation studies, we compare the results of PRMs and ORMs from the same model checkpoint.
@@
To further evaluate the alignment of PRM’s Best-of-N selections with physician preferences, nine physicians reviewed the top selections from various PRM checkpoints across 2,000 generations produced by LLaMA-Clinic. Each experiment involved at least three physicians and was conducted in a randomized and blinded manner.

### v3

@@ Abstract
Process-supervised reward models (PRMs) excel at providing step-by-step verification for large language model (LLM) outputs in domains like mathematics and coding. However, their application to fields lacking ground-truth answers, such as clinical note generation, poses significant challenges. We introduce a novel framework for training PRMs to deliver step-level reward signals for LLM-generated clinical notes. By precisely defining meaningful “steps,” injecting realistic “errors” informed by domain expertise, and leveraging LLMs to generate process supervision data at scale, we overcome previous limitations. Our PRM, built on LLaMA-3.1 8B, consistently outperforms proprietary reasoning and non-reasoning models, achieving state-of-the-art (SOTA) performance on two key evaluations: (1) distinguishing gold-standard from error-containing samples with 98.8% accuracy, and (2) selecting physician-preferred clinical notes with 56.2% accuracy. We investigate critical components for effective PRM training, including optimal loss functions and data selection strategies, and present a comprehensive physician reader study identifying predictors of downstream Best-of-N performance. Our study sheds light on unlocking the potential of PRMs for diverse generative tasks across domains. [^1] Our code is available at https://github.com/hanyin88/prm-clinic.
@@
Hanyin Wang1,2, Chufan Gao2, Qiping Xu1, Bolun Liu1, Guleid Hussein1, Hariprasad Korsapati1, Mohamad El Labban1, Kingsley Iheasirim1, Mohamed Hassan1, Gokhan Anil1, Brian Bartlett1, Jimeng Sun2,3 1 Mayo Clinic Health System, 2 School of Computing and Data Science, UIUC 3 Carle Illinois College of Medicine, UIUC wang.hanyin@mayo.edu, jimeng@illinois.edu
@@ 1 Introduction
Figure 1: Overview of study design. We train a PRM to step-by-step verify clinical notes generated by LLMs from patient-doctor dialogues. Step 1: Clinical notes are transformed into a hierarchical structure of steps, designed based on domain expertise to capture the key considerations in clinical documentation. Step 2: Gemini Pro 1.5 is utilized to generate synthetic errors from predefined categories. These errors are systematically swapped with the original steps to create negative samples. Step 3: The PRM is trained using LLaMA-3.1 8B instruct, followed by ablation studies and a physician reader study.
@@
LLMs show promise in generating clinical notes Van Veen et al. (2024), but their outputs may contain errors and misalign with physician preferences Omiye et al. (2024); Jin et al. (2024). Currently, no automated, scalable method exists to evaluate the quality of LLM-generated clinical notes, leaving manual evaluation the gold standard. Consequently, costly clinician reader studies are required to validate LLM-based products, such as ambient AI scribes Liu et al. (2024). This challenge is amplified by the rapid adoption of ambient scribing technology in the real world Tierney et al. (2025), with projections indicating that 30% of the healthcare market will use LLM-generated clinical notes by 2025 Beavins (2024).
@@
A potential solution to this challenge is to verify LLM outputs using a reward model (RM). RMs are integral to reinforcement learning for LLM post-training Ouyang et al. (2022); Stiennon et al. (2020). When employed as verifiers, RMs are categorized into two types: outcome-supervised reward models (ORMs) Cobbe et al. (2021) and, more recently, process-supervised reward models (PRMs) Lightman et al. (2023). ORMs evaluate an entire generation and assign a single reward score at the end, whereas PRMs provide reward scores at each step of the generation, offering several notable advantages. PRMs enable fine-grained verification of LLM outputs, improving explainability by pinpointing errors at their exact location. This step-wise reward signaling also facilitates inference-time scaling, such as Monte Carlo Tree Search (MCTS), and supports step-by-step reinforcement learning Wang et al. (2023b); Snell et al. (2024).
@@
However, despite significant success in mathematical and coding tasks Lightman et al. (2023); Wang et al. (2023b); Ma et al. (2023), the application of PRMs to other domain remains underexplored Beeching et al. . This gap arises from several key challenges: (1) Verification of correctness: Unlike mathematical or coding problems, which have objective ground-truth answers, verifying correctness in open-ended generative tasks is inherently more complex. (2) Process-supervised data collection: Training PRMs requires datasets with annotations for each step of the answer, which makes it difficult to scale when relying on human annotators.
To address these challenges, we introduce a novel framework for training PRMs to verify open-ended text generation in the clinical domain (Figure 1). We specifically target the rapidly growing use case of ambient scribing, where LLMs generate clinical notes from patient-doctor dialogues Barr et al. (2024). Drawing on domain expertise, we carefully design clinically meaningful “steps” and inject realistic “errors” to reflect authentic documentation practices. To scale supervision, we employ an LLM to generate process-labeled data by systematically introducing errors into high-quality reference notes. Our main contributions are as follows:
@@
    Our PRM, built on LLaMA-3.1 8B, significantly outperforms both proprietary reasoning and non-reasoning models for clinical note verification. It achieves 98.8% accuracy in detecting erroneous samples and 56.2% accuracy in selecting physician-preferred notes, setting a new SOTA in this critical domain.
@@
    We broaden the scope of PRMs beyond domains with clear ground-truth answers, introducing a generalizable framework that integrates LLMs with domain knowledge to generate process-supervised data at scale.
@@
    To support future research, we release our full dataset, including process supervision annotations and physician preference labels.
@@ 2 Background and Related Work
Recent research highlights a growing interest in process supervision of LLMs to improve their reasoning capabilities. In contrast to earlier work on ORMs Cobbe et al. (2021); Yu et al. (2024), PRMs assess the correctness of LLM outputs step-by-step Lightman et al. (2023); Uesato et al. (2022). One of the core applications of PRMs is Best-of-N sampling, where N complete solutions are generated, and the most suitable one is selected based on the PRM’s reward signal. This method has proven exceptionally effective in mathematical reasoning and code generation tasks, frequently surpassing ORMs Wang et al. (2023b); Ma et al. (2023). Furthermore, PRMs can facilitate step-level reinforcement learning and guide inference-time scaling through techniques such as MCTS Wang et al. (2023b); Ma et al. (2023); Snell et al. (2024). They are considered crucial to the success of advanced reasoning models Jiang et al. (2024); Qin et al. (2024); Wang et al. (2024b).
@@
Acquiring human-labeled annotations for each step of PRM training data poses a prohibitive scaling cost, a challenge central to our work. To address this limitation, several studies have proposed methods to automatically generate process supervision data, using techniques such as Monte Carlo estimation Wang et al. (2023b); Luo et al. (2024); Wang et al. (2024c). However, these approaches are primarily confined to the mathematical domain, where they are effective due to the unambiguous and easily verifiable nature of mathematical solutions. Our exploration of PRM is also related to prior work on evaluating step-by-step reasoning chains Hao et al. (2024), though our focus is on open-ended text generation tasks, which present unique challenges.
@@
Ambient scribe systems that transcribe patient-provider conversations into clinical notes are among the most impactful applications of LLMs in healthcare, offering substantial reductions in documentation burden Barr et al. (2024); Tierney et al. (2025). This task was explored at the 2023 ACL ClinicalNLP and CLEFImage workshops, where the best results were achieved using proprietary models Abacha et al. (2023); Yim et al. (2023a). Recent studies also show promise in fine-tuning open-source LLMs to generate expert-level clinical notes Wang et al. (2024a, 2023a); Li et al. (2024), though challenges such as inaccuracies and hallucinations remain.
@@ 3 Our Approach: Building PRMs to Verify Clinical Note Generation
In this section, we first introduce the task, followed by an explanation of the core method for constructing the process supervision dataset. We then discuss the training, usage, and evaluation methods for PRMs. An overview of our study is illustrated in Figure 1.
@@ 3.1 Problem Formation
Given a patient-doctor dialogue, an LLM can generate clinical notes comparable to those written by physicians. We aim to train PRMs to effectively verify LLM-generated notes step-by-step, enabling them to evaluate multiple candidates and select the most accurate and physician-preferred note. Guided by recent work (see details in Appendix D.3), we target the most challenging “Assessment and Plan” (A&P) section of the note.
@@ 3.2 Training Data Construction
Unlike mathematical problems, where each solution step is typically a single sentence, the concept of a “step” in an open-ended text generation task is less clear. The A&P sections of clinical notes are semi-structured, consisting of multiple problems, each encompassing several clinical reasoning narratives related to their assessment and plan (Figure 1 and 2). While various design options exist, we adopted a heuristic approach informed by physician expertise (see Section D.4 for details), as outlined below.
@@
  • Factual Inaccuracy: Errors involving incorrect information that are referenced in the conversation but not supported by its content.
@@
We tasked Gemini Pro 1.5 Team et al. (2024) with generating a pool of unique errors for each error type in each case. This was achieved through careful prompt engineering and manual inspection of the errors by our physician co-authors. Subsequently, we randomly swapped original steps in the gold-reference notes with entries from the error pools, ensuring that each error was used only once. At this stage, we introduced a fourth error type, Incompleteness, by randomly removing specific steps or entire problems from the samples.
@@
Table 1: Summary statistics of PRM-Clinic. The reported numbers are calculated solely from negative samples. Samples from the same case share the same dialogues. During training, gold-reference samples are randomly mixed with negative samples.
Step Scores
Each step is assigned a score label of “++” or “-”. The “++” label represents a correct step, and the “-” label indicates an incorrect or incomplete step. For gold-reference samples, every step receives a “++” label. For samples containing errors, the erroneous step is assigned a score label of “-”. For samples with incompleteness, we also assign the corresponding problem-level or note-level completeness score label to “-”. Lastly, for any samples containing errors, we assign a score label of “-” to the end-of-note step.
@@
A-Prefer Physician-preference ACI-BENCH test2 and test3 subsets 80 OOD Includes physician preference data used in the first round of RLHF in Wang et al. (2024a). Each case contains three notes generated by an early phase of LLaMA-Clinic.
A-Verify Gold-reference ACI-BENCH test2 and test3 subsets 80 OOD Gold-reference notes generated by an later phase of LLaMA-Clinic. Negative samples are introduced in the same manner as PRM-Clinic.
@@ 3.3 Training of PRMs
Training Details
The training corpus consists of dialogue concatenated with clinical notes, where each step in the notes is separated by a step token followed by a score label token. We used LLaMA-3.1 8B Instruct Dubey et al. (2024) as the base model, utilizing its reserved special tokens for all step and score label tokens. Further implementation details can be found in Section B.
@@ 3.4 Usage of PRMs
At inference time, we perform a single PRM forward pass over the entire clinical note to obtain PRM’s predicted score label for each step. The clinical note itself is generated by a separate LLM. We define the PRM score as the probability that a given step is correct, which is the softmax probability of the special “++” token at that step position.
@@
Figure 3: Main results. Our PRM outperforms all baseline models, including the vanilla ORM, Gemini Pro 1.5, and SOTA reasoning models, across all evaluated tasks. For A-Prefer, the task involves selecting the physician-preferred note from three candidate notes. For all other tasks, the objective is to identify the correct gold-reference sample from negative samples. For ID tasks, the notes are generated by Gemini Pro, while for OOD tasks, the notes are generated by LLaMA-Clinic. The results of the best-performing PRM are presented alongside the ORM results from the same checkpoint. ID: In-distribution. OOD: Out-of-distribution.
@@ 3.5 Evaluation of PRMs
We evaluated the reward model’s ability to select the superior note within the Best-of-N framework using two key metrics: (1) Accuracy of selecting the gold-reference note, which measures the model’s ability to identify the gold-reference note among error-containing samples (i.e., negative samples), and (2) Accuracy of selecting the physician-preferred note, which assesses its capability to identify the physician-preferred note from a set of candidate notes.
@@
We compare our PRM against several strong baselines, including LLM-as-a-judge setups using Gemini Pro 1.0 Team et al. (2023), Gemini Pro 1.5 Team et al. (2024), OpenAI’s GPT-4o Hurst et al. (2024), o1 Jaech et al. (2024), and o3-mini OpenAI (2025). Additionally, we evaluate a vanilla ORM implementation, where the training corpus lacks step-level annotations and includes only a single score for the entire clinical note. For our ablation studies, we directly compare the performance of PRMs and ORMs derived from the same model checkpoint.
@@
To identify predictors of downstream Best-of-N performance, we conducted a comprehensive reader study involving nine physicians. For each case, LLaMA-Clinic Wang et al. (2024a) generated 2,000 notes from the same patient-provider dialogue. We then used various PRM checkpoints to select the top-scoring note for physician evaluation. Each experiment was conducted in a randomized and blinded manner, involving at least three physicians. Additional details are provided in Section 4.3.
Table 3: Ablation studies on loss functions and data selection. We evaluated different strategies for cross-entropy loss computation and training data selection. The best performing model used notes-only loss and the full dataset with paraphrases. PRM and ORM results are reported from the checkpoint with the highest PRM A-Prefer performance. Numbers represent percentages of accuracy. Bold values indicate the highest PRM and ORM accuracy in each ablation study for OOD tasks. In ID tasks, multiple models achieved perfect accuracy.